OpenStudy (anonymous):

Clare has 130 songs in her music collection. another percent question i might know the answer but im not sure. Fifty-nine of the songs are rock and roll. About what percent of the songs are rock and roll? A.45% B.59% C.71% D.41%

hartnn (hartnn):

whats 59*100/130 ?

OpenStudy (anonymous):

what? i dont understand

hartnn (hartnn):

did you do it the same way ?

hartnn (hartnn):

% rock and roll = no. of rock and roll songs *100/ total no. of songs

hartnn (hartnn):

doesn't that make sense ? ^ any doubts ?

OpenStudy (anonymous):

ya. can you make it a little simpler to read i dont get it. please

OpenStudy (anonymous):

I think that he means to do this 

hartnn (hartnn):

not that,no.... and if you have 4 apples and i took 2, how many % of apples do you have ? 50%, right ? how did you get 50 ? % = 2*100/4 = 200/4 = 50, right ? similarly, for 59 rock and roll songs, out of 130 % = 59*100/130

OpenStudy (anonymous):

yess @hartnn is right lol what I did was wrong

OpenStudy (anonymous):

thanks both of you i got my answer

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

u welcome

OpenStudy (anonymous):

can one of you give me a medal please.

OpenStudy (anonymous):

thanks for the medals